OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? 1 2 3 4 5

OpenStudy (anonymous):

x+x^2=20 x^2+x-20=0 (x+5)(x-4)=0 x=-5 and x=4

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

-5 is not positive hahd

OpenStudy (anonymous):

lol